OpenStudy (anonymous):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have? 0 1 2 infinitely many

OpenStudy (anonymous):

hoowaahh i think i can help you

OpenStudy (anonymous):

Lol haha

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. slope of 1 line is 4, and has a y intercept of -3 y = 4x - 3 slope of line 2 is 4, and has y intercept of 8 y = 4x + 8 Now look at this.. if the lines have the same slope and the same y intercept, then there are infinite solutions. if the slopes are the same and the y intercepts different, then the lines are parallel with no solutions. if the slopes and the y intercepts are different, then there is 1 solution. So what do you think ?

OpenStudy (anonymous):

We dont have a y intercept yet though...hmm. We probably can get the same y intercept for the two

OpenStudy (texaschic101):

yes we do...the y intercepts are given to you in the problem. line 1, slope 4, y intercept -3 line 2, slope 4, y intercept 8

OpenStudy (anonymous):

hull up hull up hull up bruh did you put it in point slope form first

OpenStudy (anonymous):

Infinitely many then and idk ^

OpenStudy (texaschic101):

wrong...same slope and different y intercepts....parallel lines...0 solutions (because they never intersect)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

But they both have the Y thats given to us in the problem

OpenStudy (anonymous):

if they are parallel then they have no solutions because they never touch

OpenStudy (texaschic101):

if the slopes are the same, and the y intercepts different, then the lines are parallel with 0 solutions

OpenStudy (anonymous):

Yeah I know..but i wouldnt be able to tell if something is parallel by just looking at something like this ---> Y = -x + 3

OpenStudy (anonymous):

and Y = -x + 8

OpenStudy (anonymous):

if it has the same slope then you know they are parallel

OpenStudy (anonymous):

Oh...how about Y = 4X + 6, and Y = 4X - 5?

OpenStudy (anonymous):

they are parraller. again same slope

OpenStudy (texaschic101):

if you have 2 of them..you can just look at tell.. y = -x + 8 y = -x + 5 same slope, different y intercepts, 0 solutions y = -x + 8 y = -x + 8 same slope, same y intercept, infinite solutions y = -x + 8 y = 2x + 3 different slopes, different y intercepts, 1 solution

OpenStudy (anonymous):

yea tex is right

OpenStudy (anonymous):

OHHHHH!!!! If the ''-x'' is the same, and the first part looks like this: Y = -x...but the numbers are different, its the same slope and different y intercepts

OpenStudy (anonymous):

yes so parallel

OpenStudy (texaschic101):

In y = mx + b form, the number in the m position is the slope and the number in the b position is the y intercept. just compare them

OpenStudy (anonymous):

What if the signs are different? like the + and -

OpenStudy (anonymous):

they are not parallel

OpenStudy (texaschic101):

that makes a difference.. y = -x + 8 y = x + 2 different slopes, different y intercepts, 1 solution

OpenStudy (anonymous):

But on the graph they look the same...just sort-of combined plots

OpenStudy (anonymous):

no because they intersect at one point

OpenStudy (texaschic101):

you have the same slopes, but different y intercepts...parallel lines...0 solutions

OpenStudy (anonymous):

Yes!! I kinda knew it..kinda.

OpenStudy (anonymous):

just go with your gut feeling

OpenStudy (anonymous):

My gut feeling is kinda always wrong when it comes to virtual school math and graphs..

OpenStudy (anonymous):

When I know something though, I stick with it until proven wrong if I am.

OpenStudy (anonymous):

im in virtual school too bruh

OpenStudy (texaschic101):

just remember, if you want to find slope and y intercept, put your equation in y = mx + b form first

OpenStudy (anonymous):

thanks bruh. ican teach you how to nae nae

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

im a girl o-o

OpenStudy (anonymous):

oh sista from anotha mista

OpenStudy (anonymous):

hull up hull up hull up HOOOOWWWWWAAAHHH

OpenStudy (anonymous):

what kinda thirsties do you slap?

OpenStudy (anonymous):

how did you change your picture too. thats funny with Obama